### v2

Abstract
Despite recent advances in multimodal large language models (MLLMs), their development has predominantly focused on English- and western-centric datasets and tasks, leaving most of the world’s languages and diverse cultural contexts underrepresented. This paper introduces Pangea, a multilingual multimodal LLM trained on PangeaIns, a diverse 6M instruction dataset spanning 39 languages. PangeaIns features: 1) high-quality English instructions, 2) carefully machine-translated instructions, and 3) culturally relevant multimodal tasks to ensure cross-cultural coverage. To rigorously assess models’ capabilities, we introduce PangeaBench, a holistic evaluation suite encompassing 14 datasets covering 47 languages. Results show that Pangea significantly outperforms existing open-source models in multilingual settings and diverse cultural contexts. Ablation studies further reveal the importance of English data proportions, language popularity, and the number of multimodal training samples on overall performance. We fully open-source our data, code, and trained checkpoints, to facilitate the development of inclusive and robust multilingual MLLMs, promoting equity and accessibility across a broader linguistic and cultural spectrum.
Figure 1: Overview of the aggregate performance of various multimodal LLMs on PangeaBench. Our Pangea-7B demonstrates comparable performance to SoTA open-source models in English settings, while significantly outperforming them in multilingual scenarios.
1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023b; 2024; Dubey et al., 2024; Deitke et al., 2024; OpenAI, 2024b; Team et al., 2023) have demonstrated impressive capabilities in tasks such as image captioning, visual question answering, and complex reasoning (Yue et al., 2024a; b). Despite this rapid progress in their reasoning capabilities, a critical flaw persists: the overwhelming focus on English- and western-centric training datasets and evaluation benchmarks (Liu et al., 2021; Song et al., 2023).
This homogeneity results in a lack of representation for the vast majority of the world’s languages and diverse cultural contexts (Yu et al., 2022). Consequently, models predominately trained on such data exhibit: (a) diminished performance in multilingual settings (Blasi et al., 2022) with poor tokenization further leading to higher inference costs (Ahia et al., 2023); (b) generate outputs misaligned with the socio-cultural norms of underrepresented languages (AlKhamissi et al., 2024); and (c) lack the ability to recognize objects from geographically diverse regions (Ramaswamy et al., 2024) or rare objects belonging to the long-tail (Gupta et al., 2019). With the increased adoption of these models into real-world applications across the globe, there’s an urgent need to develop multilingual MLLMs that equitably serve a diverse set of users. Few efforts have been made to develop multilingual MLLMs (Geigle et al., 2024b; Rasheed et al., 2025), however, their performance still exhibits inequalities across languages and lacks evaluation of cultural understanding.
In this paper, we address how to train and evaluate multilingual MLLMs that are culturally inclusive, using limited open-source resources, tackling four major challenges (Yu et al., 2022): 1) Data scarcity: high-quality multilingual multimodal data is scarce, especially in low-resource languages, which makes it difficult to create large-scale training data; 2) Cultural nuances: visual interpretations are context-dependent and vary across cultures (Ramaswamy et al., 2023; Khanuja et al., 2024); 3) Catastrophic forgetting: training on many languages or modalities often results in suboptimal performance on some subsets and require careful balancing; 4) Evaluation complexity: developing an evaluation suite that accurately measures performance across languages and cultures requires substantial resources and expertise.
To tackle these challenges, we introduce Pangea, an open-sourced multilingual MLLM designed to bridge linguistic and cultural gaps in visual understanding tasks. Pangea is trained on PangeaIns (Figure 2), a high-quality multilingual multimodal instruction tuning dataset comprising 6 million samples in 39 typologically diverse languages. PangeaIns combines existing open-source resources with newly created instructions focused on multicultural understanding. We curate high-quality English instructions, carefully translate them, and adapt them for multilingual contexts. To address Western-centric biases in visual representations, we source images from LAION-Multi (Schuhmann et al., 2022), which includes images from various countries and captions in multiple languages. However, LAION-Multi contains many images that are not culturally representative of the country’s speaking population. Additionally, the associated alt text is often short, noisy, and lacks sufficient detail. To combat these issues, we develop a multi-cultural multilingual multimodal instruction generation pipeline. This pipeline leverages an LLM (Dubey et al., 2024) to score and filter images based on cultural informativeness. We then enhance the remaining data by generating detailed descriptions and creating complex instructions that combine culturally relevant tasks with general multilingual scenarios. This approach improves the model’s cultural understanding while maintaining robust multilingual performance.
To evaluate Pangea’s capabilities, we present PangeaBench, a comprehensive multilingual and multimodal evaluation suite comprising five multimodal and three text-based tasks across 14 datasets in 47 languages. PangeaBench assesses MLLMs’ performance on open-domain multimodal chat, image captioning, cultural understanding, multimodal reasoning, and text-only tasks including question answering and complex math reasoning. A key highlight of PangeaBench is the introduction of xChat, a human-crafted benchmark designed to evaluate open-ended, information-seeking multimodal conversations. xChat employs a fine-grained evaluation pipeline where human annotators annotate both reference answers and scoring rubrics for each query. An LLM then uses these rubrics to score the model’s predictions on a 1-5 scale. This approach offers a more precise assessment of MLLM performance, addressing limitations of coarse LLM-as-Judge methods (Zheng et al., 2023). Additionally, we introduce xMMMU, a translated version of MMMU (Yue et al., 2024a), testing college-level multimodal reasoning across seven languages. Together, these components provide a robust, nuanced evaluation of Pangea’s cross-lingual and cross-cultural capabilities.
Our results demonstrate Pangea’s abilities in both English and multilingual scenarios, significantly outperforming existing open-source MLLMs across PangeaBench datasets, surpassing the best open MLLMs by 0.4 points on English tasks and 10.9 points on multilingual tasks on average. Notably, Pangea excels in multilingual and multi-cultural understanding, as evidenced by its performance on xChat, CVQA, and MaRVL benchmarks. Pangea also matches or outperforms state-of-the art proprietary LLMs, namely Gemini-1.5-Pro and GPT4o, on several tasks such as XGQA. However, some performance gaps remain in multimodal chat and complex reasoning, shedding light on the need for further improvements in open multimodal LLMs. We discuss key insights from Pangea’s training, including the scaling effect of instructions, the role of English data, and the impact of language-specific training proportions on performance. Additionally, we explore preliminary methods to improve multilingual OCR. We fully open-source PangeaIns, PangeaBench, Pangea models, and code, to advance culturally inclusive MLLMs across diverse languages.
Figure 2: Statistics of our PangeaIns, which comprises 6M multimodal instructions in 39 languages. The distribution of multilingual training samples shows the percentage of instances for each language among the multilingual instructions. PangeaIns includes general instructions, document and chart question answering, captioning, domain-specific, culturally relevant, and text-only instructions. To address the data scarcity of multilingual multimodal instructions, we employ three strategies: 1) translating high-quality English instructions into multiple languages, filtering culturally relevant images; 2) generating culturally aware instructions, and 3) collecting and re-purposing existing open-source multimodal datasets.
2 PangeaIns: Multilingual Multimodal Instruction Tuning
Creating a truly multilingual, multicultural MLLM presents unique challenges. We developed PangeaIns, a diverse and high-quality dataset specifically designed for instruction tuning. Comprising 6 million samples in 39 languages, PangeaIns was curated with a focus on linguistic and cultural diversity. We implemented three key strategies to ensure comprehensive coverage, each addressing the specific hurdles encountered in multilingual multimodal learning. Figure 2 shows the details of our PangeaIns’s distribution.
2.1 Machine Translated Instructions
The scarcity of human-annotated multilingual multimodal training datasets presents a significant challenge. To address this, we primarily adopt machine translation as a practical and scalable solution to extend our dataset beyond English. While human annotation is ideal, it is resource-intensive and impractical for covering a wide range of languages.
Constructing a High-quality Pool of English Instructions from Existing Sources. We first collect a high-quality set of English multimodal instructions, which serve as the foundation for translation into other languages. These instructions span a wide range of visual understanding tasks, including general visual instructions and conversations (Tong et al., 2024; Liu et al., 2024), visual reasoning, captioning, and chart question answering (Masry et al., 2022). Besides, we also added text-only high-quality English instructions, covering general instructions (Teknium, 2023), code (Zheng et al., 2024), and math (Li et al., 2024c). Figure 2 shows the statistics of our translated datasets. By leveraging existing English instructions, we ensured comprehensive coverage of visual interpretation and text instruction following tasks in English, preparing a pool of high-quality data for translation.
Translation Model Selection. To expand the English instructions to other languages, we initially experimented with strong open-source machine translation models, such as NLLB-3B (NLLB Team, 2024). However, we found that these models struggled with complex instruction-following scenarios and context-switching tasks, particularly in specialized domains like code generation and mathematical reasoning. For example, in code-related tasks, the model failed to recognize and correctly translate programming language keywords, significantly reducing the quality of the instructions. Based on these limitations, we shifted to using the proprietary Gemini 1.5 Pro model, which shows slightly better performance in small-scale human evaluations compared with GPT4o.
Post-Processing Translated Data. Even with high-quality translations, inconsistencies arose. To resolve issues such as mismatched conversation turns or missing candidates in multiple-choice questions, we developed a post-processing pipeline. This pipeline automatically corrected these errors or directly dropped the examples, ensuring that all translated instructions remained consistent.
Overall, Gemini 1.5 Pro’s translation seems satisfactory, providing a fast, cost-effective alternative to human annotation, especially for scaling across languages. However, we acknowledge that machine translation still has limitations, particularly in handling nuanced contexts and cultural subtleties.
Figure 3: Overview of multicultural understanding instructions data generation pipeline.
2.2 Multicultural Understanding Instructions
While machine translation enables scaling across multiple languages, data translated from English is still Anglo-centric in coverage of cultural concepts (Yu et al., 2022). To address this, we developed a pipeline focused on creating instructions for multicultural understanding. Both visual and textual elements can convey deep cultural significance, and our goal is to design a dataset that allows models to not only recognize these nuances but also respond appropriately across various cultural contexts. The pipeline of creating multicultural understanding instructions is shown in Figure 3.
Curation of Culturally Diverse Images. To ensure that our dataset captures a wide array of cultural contexts, we began by sampling 10 million images from the LAION-Multi dataset (Schuhmann et al., 2022), which includes images and short alt texts from diverse languages and regions. A filtering process was proposed to guarantee both the quality and cultural relevance of the images.
  • Heuristic Filtering: We implemented automatic filtering based on several key criteria: Image Size, Aspect Ratio, Text Length, NSFW content, Offensive Text, Deduplication, and CLIP Score (used to assess the alignment between the image and its textual description). This helped remove low-quality or inappropriate images and ensured the remaining dataset adhered to quality standards.
  • LLM Scoring: To further refine the dataset, we employed the Llama-3.1-8B-Instruct model (Dubey et al., 2024) to evaluate the quality, subjects, and cultural relevance of the accompanying text descriptions (alt text) for each image. The model was instructed to perform the following tasks: 1) Evaluate Text Quality: The alt text was rated on a scale from 1 to 5 based on how well it described the corresponding image, assuming the model could not access the image itself. Alt text scoring below 4 was removed. 2) Subject Classification: The model assigned a subject or category to the alt text based on its content. 3) Country/Region Classification: The model determined whether the alt text was closely related to a specific country’s culture. Images classified as “no specific country” (approximately 60% of the dataset) were excluded to ensure we focused on culturally identifiable content. The full LLM scoring prompt is included in Appendix B.
  • Avoiding Overrepresentation: To maintain a balanced representation, we downsampled images from frequently occurring subjects, such as objects, materials, and clothing, to avoid skewing the dataset toward specific topics or regions. Then, we conducted an accessibility check, removing 30% of the remaining samples due to image download or other issues. Ultimately, we curated a final set of 1M high-quality, culturally specific images, forming the foundation of PangeaIns.
Captioning of Multicultural Images with Different Languages. To provide context and enhance the models’ ability to interpret the images accurately, we regenerated a more detailed caption using Gemini 1.5 Pro based on high-quality original text. Each image was accompanied by a caption written in the language corresponding to its cultural origin. This step was crucial as writing captions in the native language of the image’s cultural context adds a layer of authenticity and helps the model learn language-specific nuances.
Generating Multilingual and Cross-Cultural Instructions. For each image, we generated captions in the corresponding native language using Gemini 1.5 Pro. However, our approach was not just about using a capable model. The filtered high-quality alt text played a critical role in enriching the data, as it often contained culturally specific and contextually important information that would otherwise be absent from the images alone. For example, in the Appendix Figure 8, with high-quality alt text, a model can incorporate details such as “President and CEO of The Walt Disney Company standing in front of a model of Shanghai Disneyland,” which adds significant context that may not be immediately evident from the image. This additional layer of information helps the model generate captions that better capture the cultural and contextual nuances.
After recaptioning, we generated multilingual instructions based on the detailed captions with Gemini 1.5 Pro. Instead of only prompting the model to generate random instructions, we did a careful prompt engineering where we first came up with 13 task types (e.g., Information Seeking, Coding & Debugging, Critical Reasoning, Cultural Interpretation, etc.). Then for each image, up to two QA pairs were created, representing different instruction types to ensure a diverse set of interactions. This approach ensures that the model not only recognizes these visual elements but also responds appropriately across varied linguistic and different instruction contexts. The captioning and instruction generation prompts are included in Appendix B.
2.3 Curating Existing Multilingual Instructions
To further enrich PangeaIns, we conducted an extensive survey of the available multilingual multimodal literature and datasets, including those hosted on HuggingFace. As a result, we incorporated several high-quality, open-source datasets into our PangeaIns mixture. These include Chinese ALLaVA-4V (Chen et al., 2024), Viet Document and OCR QA (Doan et al., 2024), Llava Chinese (LinkSoul-AI, 2023), Llava Medical Chinese Instruction (BUAA, 2023), LLaVA-Japanese-Instruct (Toshi456, 2023), MTVQA (Tang et al., 2024), Japanese STAIR Captions (Yoshikawa et al., 2017), Russian GQA (Belopolskih & Spirin, 2024), French Doc-VQA (Sonagu & Sola, 2024), and French Table-VQA (Agonnoude & Delestre, 2024). Each of these datasets brings unique linguistic and cultural perspectives to the mix, covering a wide range of languages and task types.
2.4 Dataset Statistics
By combining these three methods, we created PangeaIns as a comprehensive dataset that addresses the major challenges in building multilingual multimodal models: data scarcity, linguistic diversity, and cultural nuance. The dataset’s balanced distribution across languages and tasks supports the development of more sophisticated LLMs that can handle complex visual and textual content in a multilingual, multicultural context.
Language and Task Distribution: PangeaIns features an extensive and balanced distribution of languages, tasks, and cultural contexts (as shown in Figure 2). We empirically keep the final language ratio of English to Multilingual as 40%:60% as we found a significant portion of English data plays an important role in cross-lingual transfer. See more discussions about the ratio in section 5 and Figure 6. The inclusion of diverse multimodal instructions ensures that the model develops a deeper understanding of varied linguistic and cultural environments. Examples of training samples from different languages and categories are provided in Appendix H.
The comprehensive nature of PangeaIns lays a solid foundation for training Pangea, enabling it to become a truly multilingual, multicultural multimodal LLM, capable of understanding and interacting effectively with users from diverse linguistic and cultural backgrounds.
Category Tasks Datasets Forms Size Languages Metric
Multimodal Multimodal Chat xChatBench Long 400 zh,en,hi,id,ja,rw,ko,es LLM-as-Judge
M-LlavaBench Long 600 ar,bn,zh,fr,hi,ja,ru,es,ur,en LLM-as-Judge
Captioning XM100 Long 3.6K 36 languages ROUGE-L
Cultural Understanding CVQA MC 21K en,zh,ko,mn,ja,id,jv,min,su Accuracy
MaRVL Short 6K id,sw,ta,tr,zh Accuracy
Multilingual VQA xGQA Short 77K en,de,pt,ru,id,bn,ko,zh Accuracy
MaXM MC 2K hi,th,zh,fr,en,iw,ro Accuracy
Reasoning (Multi-subject) xMMMU Short/MC 3K en,ar,fr,hi,id,ja,pt Accuracy
M3Exam MC 3K en,zh,it,pt,vi,th,af Accuracy
Text-only QA TyDiQA Short 5.1K ar,ru,bn,te,fi,sw,ko,id,en Accuracy
Translation FLORES-Sub Long 18K ar,en,fr,de,hi,id,iw,ja,pt,ro,tr ChrF
Reasoning (Multi-subject, Commonsense, Math) MMMLU MC 197K ar,bn,de,es,fr,hi,id,it,ja,ko,pt,sw,yo,zh Accuracy
XStoryCloze MC 21K en,ar,es,eu,hi,id,my,ru,sw,te,zh Accuracy
MGSM Open 3K bn,de,en,es,fr,ja,ru,sw,te,th,zh Accuracy
Figure 4: Overview of PangeaBench, which contains 5 multimodal and 3 text tasks covering 14 datasets (including two newly curated xChatBench and xMMMU datasets). The table provides details about the datasets, while the figure shows evaluation examples from five different multimodal eval tasks in our PangeaBench.
3 PangeaBench: Evaluation of Multilingual Multimodal Models
3.1 Overview of PangeaBench
To assess the capabilities of Pangea across a variety of languages, cultures, and task types, we have developed PangeaBench, a comprehensive multilingual and multimodal evaluation suite. PangeaBench integrates diverse benchmarks that encompass both multimodal and text-only tasks, enabling a holistic evaluation of Pangea’s performance in cross-lingual, cross-cultural, and multimodal contexts. Each task within PangeaBench is designed to probe specific aspects of Pangea’s proficiency, ensuring robust testing across a wide range of scenarios. All tasks in PangeaBench are evaluated under a zero-shot setting.
3.2 Multimodal Tasks
The multimodal tasks in PangeaBench are categorized as follows: Multimodal Chat, Captioning, Cultural Understanding, Multilingual Visual Question Answering (VQA), and Multi-Subject Reasoning. We incorporate these tasks in PangeaBench to ensure comprehensive testing of Pangea’s multimodal capabilities. The overview and examples of PangeaBench from each task are shown in Figure 4.
Multimodal Chat. This task tests the model’s ability to engage in natural and dynamic real-world conversations involving both text and images. In this category, Multilingual LlavaBench (Rasheed et al., 2025) (M-LlavaBench for short) stands as the only benchmark for evaluating multilingual long-form generation capabilities from MLLMs. Moreover, following the evaluation pipeline from Zheng et al. (2023) and Liu et al. (2023b), M-LlavaBench uses a coarse-grained evaluation criteria (for example, “Please rate the helpfulness, relevance, accuracy, level of details of their responses.”). Previous works have shown that employing such coarse-grained evaluation criteria might lead to automatic evaluation results that diverge from how humans would evaluate them (Ye et al., 2023; Kim et al., 2023; Lee et al., 2024; Kim et al., 2024a; b). To assess our baselines based on a more accurate evaluation pipeline with fine-grained evaluation criteria on diverse scenarios, we additionally annotate a new multilingual multimodal generation benchmark called the xChatBench, included in the multimodal chat category of PangeaBench. A more detailed explanation of the annotation process of xChatBench is included in Appendix E.
Captioning. The XM3600 (Thapliyal et al., 2022) dataset was developed to evaluate models’ capability in multilingual image captioning. This dataset contains images paired with captions in 36 different languages. The original datasets often contain similar images and captions. To address this, we clustered the images based on their captions and manually selected 100 representative images from these clusters. This approach enhances the diversity of the dataset while also accelerating the evaluation process. We call the dataset XM100.
Cultural Understanding. To assess the model’s ability to reason about and understand culturally diverse visual content, we use the CVQA (Romero et al., 2024) and MaRVL (Liu et al., 2021) datasets. These datasets are designed to test the model’s performance in reasoning tasks involving culturally relevant imagery and concepts across multiple languages.
Multilingual VQA. This task measures the model’s proficiency in answering questions about images across multiple languages. The xGQA (Pfeiffer et al., 2022b) and MaXM (Changpinyo et al., 2022) datasets provide a diverse range of visual question-answering challenges in several languages and scripts, addressing cross-lingual visual understanding.
Multi-Subject Reasoning. The xMMMU and M3Exam (Zhang et al., 2023) datasets are used to evaluate the model’s reasoning abilities across different academic subjects. xMMMU is a machine-translated version of MMMU validation questions, which focuses on multimodal reasoning in multiple subjects. We randomly sample 300 questions from MMMU (Yue et al., 2024a) validation set and employ GPT-4o for the six languages translation. M3Exam challenges the model with complex, real-world educational questions requiring both textual and visual comprehension. Further details on how we ensure the quality of translations for xMMMU, as well as a detailed description of other datasets, can be found in Appendix D.
3.3 Text-Only Multilingual Datasets
While multimodal tasks are critical for evaluating the holistic capabilities of models like Pangea, text-only multilingual tasks provide an equally essential dimension to assess. Most existing multimodal evaluations tend to overlook the importance of text-only evaluation, especially across diverse languages. Including text-only tasks in PangeaBench allows us to examine whether the model can perform well in scenarios that require deep linguistic understanding without the aid of visual context, highlighting its performance as a foundation model. We include three tasks QA, Translation, and Reasoning covering five datasets for the text-only evaluations in PangeaBench.
Specifically, we include TydiQA (Clark et al., 2020) to test the model’s ability to answer questions across 11 typologically diverse languages. We adopt the FLORES (NLLB Team, 2024) dataset to assess machine translation performance across multiple languages. We sample 11 languages (denoted as FLORES-Sub). We use MMMLU (OpenAI, 2024a), a human-translated version of MMLU to test the general language understanding of models. We use XStoryCloze (Lin et al., 2021) and MGSM (Shi et al., 2022) to test the model’s commonsense and mathematical reasoning ability in multilingual contexts respectively.
4 Experiments
4.1 Experimental Setup
Stages Pretraining Finetuning
Training Data
Dataset LLaVA LCS-558K PangeaIns
#Samples 558K 6M
Model
Trainable Projector (20M) Full Model (8B)
Training
Batch Size 128 128
LR: $\psi_{\text{vision}}$ 1 $\times$ $10^{-3}$ 2 $\times$ $10^{-6}$
LR: $\{\theta_{\text{proj}},\phi_{\text{LLM}}\}$ 1 $\times$ $10^{-3}$ 2 $\times$ $10^{-5}$
Epoch 1 1
GPU Hours (H100) 32 1344
Table 1: Pangea’s training configurations.
We train Pangea on PangeaIns, our multilingual multimodal dataset comprising 6 million samples across 39 languages. The model is based on LLaVA-Next architecture (Liu et al., 2024) with Qwen2-7B-Instruct (Yang et al., 2024) as the language model backbone and clip-vit-large-patch14-336 (Radford et al., 2021) as the vision encoder. The training consists of two stages. First, we pretrain the vision-language connector that aligns the outputs of clip-vit-large-patch14-336 to Qwen2-7B-Instruct, with the LLaVA LCS-558K [^1] https://huggingface.co/datasets/liuhaotian/LLaVA-Pretrain (Liu et al., 2023b; a). Then, we perform finetuning on PangeaIns, where we employ a learning rate of 2e-5, a batch size of 512, coupled with a cosine decay schedule with 0.03 warmup steps. We pretrain and finetune the model for 1 epoch, where pretraining took 4 hours with 8 H100 (32 GPU hours), and finetuning took 168 hours with 8 H100 (1344 GPU hours).
For evaluation, we compare Pangea against several state-of-the-art open source baselines, including English-centric models Llava-1.5-7B (Liu et al., 2023a), Llava-Next-7B (Liu et al., 2024), Phi-3.5-Vision (Abdin et al., 2024), Cambrian-8B (Tong et al., 2024), Llava-OV-7B (Li et al., 2024b), Molmo-7B-D (Deitke et al., 2024) Llama3.2-11B (Dubey et al., 2024) and multilingual models PaliGemma-3B (Beyer et al., 2024), PALO-7B (Rasheed et al., 2025), mBLIP mT0-XL and mBLIP BLOOMZ (Geigle et al., 2024b). We also consider two text-only LLMs baselines Vicuna-1.5-7B (Zheng et al., 2023) and Qwen2-7B-Instruct (Yang et al., 2024), which are the backbones of Llava-Next and our Pangea respectively. We integrate our multimodal tasks in PangeaBench into lmms-eval (Li et al., 2024a), a multimodal evaluation package that supports many English multimodal benchmarks. We use lm-evaluation-harness (Biderman et al., 2024) to evaluate text-only tasks. We follow the original paper for their best models’ prompts in different tasks, and mostly reproduce their original numbers on datasets reported in the original papers.
Models AVG (all) Multimodal Chat Cultural Understanding
xChatBench M-LlavaBench CVQA MaRVL
en mul en mul en mul en mul en mul
Gemini-1.5-Pro 67.1 62.5 67.0 54.4 103.4 106.6 75.9 75.7 76.4 72.0
GPT4o 68.6 64.6 71.0 64.4 104.6 100.4 79.1 79.4 81.4 82.1
Llava-1.5-7B 45.4 28.4 28.5 11.8 66.1 40.8 48.9 36.5 56.2 53.7
Llava-Next-7B 51.1 32.7 40.5 18.9 78.9 50.7 55.7 42.6 62.8 50.9
Phi-3.5-Vision 54.0 35.0 38.5 13.2 70.8 58.0 56.3 42.3 72.1 56.5
[…]
Molmo-7B-D 55.4 34.1 49.5 21.1 95.9 13.8 59.4 48.3 65.3 54.9
[…]
PaliGemma-3B 37.3 25.8 6.0 3.5 32.1 31.9 52.9 42.9 56.5 52.2
PALO-7B 46.3 32.2 27.0 11.8 68.9 71.2 50.9 39.2 63.3 54.2
[…]
mBLIP BLOOMZ 36.1 30.0 4.0 1.6 43.5 41.0 44.9 36.9 62.3 58.6
[…]
$\Delta$ over SoTA Open +0.4 +10.9 -3.5 +7.3 -11.7 +18.3 -5.8 -4.2 +11.6 +12.3
Models Captioning Short VQA Multi-subject Reasoning
XM100 xGQA MaXM xMMMU M3Exam
en mul en mul en mul en mul en mul
Gemini-1.5-Pro 27.6 19.1 54.2 48.7 56.4 63.5 65.8 57.7 77.4 64.7
GPT4o 27.7 19.1 55.8 51.0 60.7 65.4 69.1 58.3 68.0 61.0
Llava-1.5-7B 28.6 1.1 62.0 30.6 49.8 20.4 36.2 31.5 32.3 29
[…]
Cambrian-8B 20.6 9.9 64.6 39.8 55.3 28.7 41.8 33.2 34.7 33.4
[…]
Molmo-7B-D 22.1 9.1 51.5 43.0 52.9 37.5 44.5 40.4 57.1 39.1
[…]
PaliGemma-3B 18.7 0.8 59.7 30.5 47.9 19.9 26.3 25.2 36.0 25.6
[…]
mBLIP mT0-XL 31.9 3.1 44.2 39.9 44.7 36.8 29.3 30.4 22.8 25
[…]
$\Delta$ over Best Open Model -0.2 +3.9 -0.1 +12.0 0.0 +9.4 -0.8 +2.3 +1.0 -3.7
[…]
4.2 Multilingual Multimodal Results
The results in Table 2 provide clear insights into the strengths and remaining challenges of Pangea-7B in multilingual and multimodal tasks. Key observations from the evaluation include:
Superior English and Multilingual Performance: Pangea-7B outperforms existing open-source models across both English and multilingual tasks. While concurrent multimodal models such as Molmo (Deitke et al., 2024) or Llama 3.2 show strong performance on English datasets, they struggle in multilingual evaluation settings. Particularly in multilingual subsets like xChatBench, M-LlavaBench, and MaRVL, it has achieved substantial gains, highlighting its effectiveness in both cross-lingual and cross-cultural contexts.
Balanced Cross-Language Capabilities: Unlike many models that exhibit a significant drop in performance when moving from English to multilingual tasks, Pangea-7B is relatively consistent. For instance, in Multimodal Chat tasks, the performance gap between English and multilingual remains relatively small, indicating its ability to handle multiple languages effectively.
Challenges Compared to Proprietary Models: While Pangea-7B leads in open-source models, some gaps remain when compared to closed-source models like GPT4o. Additionally, though Pangea-7B narrows the gap between English and multilingual performance, there is still room for improvement in fully closing this divide across all tasks.
4.3 Multilingual Text-only Results
We further evaluate our model in text-only scenarios in Table 3. Interesting findings include:
Best Text Performance Among Multimodal LLMs: Pangea-7B demonstrates the strongest performance among all multimodal LLMs in the text-only tasks consistently outperforming baselines like Llava-Next-7B. This highlights that, despite being trained as a multimodal model, Pangea-7B maintains superior text understanding and reasoning capabilities compared to other MLLMs.
Maintained Performance from its Text Backbone. Pangea-7B generally maintains or sees slight drops in performance on most text-only benchmarks compared with its text backbone Qwen2-7B-Instruct. Notably, the model shows a significant improvement in MGSM. This improvement is directly attributable to the inclusion of math-related instructions in PangeaIns, which enhances the model’s capability to handle complex multilingual reasoning and mathematical tasks.
Models AVG (all) FLORES-Sub TyDiQA XStoryCloze MGSM MMMLU
en mul x→en en→x en mul en mul en mul en mul
Vicuna-1.5-7B 52.1 38.7 55.6 42.4 59.7 52.7 78.1 57.4 17.6 6.4 49.5 34.7
[…]
Llava-1.5-7B 53.1 39.0 54.7 41.5 66.8 52.8 79.1 57.6 14.8 7.6 50.2 35.7
Llava-Next-7B 54.0 38.9 54.8 41.4 68.3 52.1 79.1 57.1 15.6 7.5 52.1 36.5
Phi-3.5-Vision 60.7 41.7 28.5 32.5 75.9 51.3 77.9 54.8 59.2 33.1 62.0 36.7
PALO-7B 52.0 37.5 52.9 40.4 69.4 50.8 77.4 57.2 13.6 5.8 46.7 33.4
[…]
Table 3: Overall performance on text-only multilingual benchmarks in PangeaBench.
5 Discussion
Finally, we explore the implications of our findings and their potential impact on future developments in the field. We examine the scaling effects of instruction quantity, the persistent role of English data, the relationship between training sample proportions and performance, qualitative examples of model behavior, and challenges in multilingual OCR. Through this discussion, we aim to provide a comprehensive understanding of our model and chart a course for future advancements.
Scaling Effect of Number of Instructions. Understanding how the quantity of instructions affects model performance is crucial for optimizing training strategies and resource allocation. Figure 5 reveals a clear scaling effect related to the number of instructions used during training. Performance improvements were consistent as we increased the number of multilingual instructions in PangeaIns, for both English and multilingual performance. This demonstrates the necessity of scaling multilingual multimodal instruction tuning.
Figure 5: Scaling effect of training samples on English and multilingual scores across datasets.
Role of English Data. In multilingual scenarios, English data plays a pivotal role in cross-lingual transfer. To investigate this, we sampled 500K examples from the translated data described in subsection 2.1, ensuring a consistent data distribution. We varied the ratio of English data while keeping the total number of training samples fixed at 500K. For the 17 multilingual languages in the translated subset, we evenly distributed the number of samples across languages.
Figure 6: Impact of English training data proportion on English vs. multilingual performance.
As shown in Figure 6 , English performance generally improves as the percentage of English data increases. More surprisingly, using no English data (full multilingual data) results in relatively lower multilingual performance. As we introduce more English data, multilingual performance improves, peaking at 38.7% with 40% English. However, performance drops sharply when English data reaches 100%. This suggests that English data aids cross-lingual transfer, however, over-reliance on it harms multilingual performance.
How does the proportion of training samples in a language affect downstream performance? An interesting question to ask is whether the downstream task performance is correlated with the number of training samples. Our analysis in Figure 7 revealed a nuanced relationship between training sample proportion and downstream performance. While there is a general positive correlation, the impact varies significantly across languages and tasks. For widely spoken languages with rich linguistic resources, we observed a near-linear relationship. However, for low-resource languages, even a small increase in proportion yielded disproportionately large performance gains. Interestingly, we also noted instances of positive transfer between typologically similar languages. These findings suggest that strategic allocation of training samples, considering both language prevalence and linguistic similarities, can optimize overall model performance.
Figure 7: The relationship between training sample size (relative to English) and performance (relative to English) of different languages across four datasets.
Multimodal Chat in the Wild. One important application of MLLMs is to answer users’ queries in the wild. Here, we show the outputs of Pangea for the multimodal chat queries from our xChatBench. As shown in Appendix Figure 9, 10, 11, 12, 13, 14, Pangea successfully interprets the figures in different tasks and generates fluent and readable in certain languages. These qualitative examples further demonstrate the remarkable visual understanding ability of Pangea in multilingual contexts. On the other hand, we also identified a few bad cases shown in Figure 15, 16. Despite generating relevant responses to the queries, the model does not capture the key details of the images due to the lack of knowledge, which points out potential improvement directions in the future.
Preliminary Explorations of Multilingual OCR. Multilingual OCR emerged as a particularly challenging aspect of Pangea’s functionality. We made efforts to enhance its multilingual OCR capabilities. Further details on exploring multilingual OCR are available in Appendix J.
6 Conclusion
In this paper, we introduced Pangea, a multilingual MLLM designed to bridge linguistic and cultural gaps in visual understanding tasks. By leveraging PangeaIns, our newly curated 6M multilingual multimodal instruction data samples, we demonstrated significant improvements in cross-lingual and cross-cultural understanding across 39 typologically diverse languages. Our comprehensive evaluation using PangeaBench revealed Pangea’s superior performance compared to existing open-source models. We also highlight ongoing challenges in areas such as low-resource language support and multilingual OCR. We fully open-source Pangea, PangeaIns, and PangeaBench to facilitate future research to build open and inclusive MLLMs.
Appendix A Related Work
Visual Instruction Tuning.
Visual instruction tuning is a key technique for enhancing multimodal large language models by aligning visual inputs with textual instructions to improve understanding and generation tasks Liu et al. (2023b). Traditionally, these instructions are built using English-language data from visual question answering and other datasets Liu et al. (2023b; a); Xu et al. (2023); Liu et al. (2024); Tong et al. (2024); Beyer et al. (2024); Zhu et al. (2024); Dai et al. (2023); Tong et al. (2024); Li et al. (2024d). Researchers often supplement this with synthetic instruction tuning data, generating large volumes of instructional pairs to possibly cover multiple languages too Geigle et al. (2024b); Li et al. (2023; 2024d). However, these instruction-tuning datasets have mostly been task-focused and lack conversational capabilities. Further, while translation gives lends to multilingual capabilities, the data remains to be culturally homogeneous. By curating multilingual and multicultural instruction tuning data across various task types, our model is designed to intuitively understand and engage with users from diverse demographics.
Dataset # Languages # of Instances Multicultural # of Task Types Open-Sourced
MultiInstruct (Xu et al., 2023) 1 $\sim$235.0K 310
MiniGPT4 (Zhu et al., 2024) 1 5.0K 149 $\checkmark$
LLAVA (Liu et al., 2023b) 1 1.2M $>$100K $\checkmark$
InstructBLIP (Dai et al., 2023) 1 $\sim$1.6M $>$100K
M3IT (Li et al., 2023) 80 2.4M 400 $\checkmark$
mBLIP (Geigle et al., 2024a) 95 5.1M 68 $\checkmark$
PALO (Rasheed et al., 2025) 10 2.1M 22 $\checkmark$
Cambrian (Tong et al., 2024) 1 7.1M $>$1M $\checkmark$
PangeaIns (Ours) 39 6.2M $\checkmark$ $>$1M $\checkmark$
Table 4: Comparison of datasets in terms of number of languages, number of instances, whether the dataset is multicultural, number of task types, and open-sourced.
Multilingual Multimodal LLMs.
Multilingual MLLMs have evolved from dual-encoder-based models, only capable of understanding and reasoning Ni et al. (2021); Zeng et al. (2022); Jain et al. (2021), to encoder-decoder models capable of multilingual text generation as well Shan et al. (2022); Chen et al. (2023); Geigle et al. (2024b). Despite their advancements, these models have remained focused on conventional tasks such as VQA and image captioning. Moreover, most efforts have centered around training with multilingual text, while little attention has been given to curating culturally diverse image datasets. Even for text, despite the focus on multilinguality, few attempts have been made to reflect cultural diversity in instructions and captions. As a result, these models tend to reflect a Western-centric bias. By selecting culturally diverse images from LAION and intentionally integrating this diversity into our instructions and captions, our model aims to serve a wide range of users in an inclusive and equitable manner.
Appendix B Prompts used in the data construction
In this appendix, we will list the detailed prompts we used when constructing cultural understanding instruction tuning data described in subsection 2.2.
Cultural Images LLM Scoring Prompt You are given an [Alt Text] associated with an image from the web. [Alt Text]: {Alt Text} Your goal is to: 1. Evaluate Text Quality: Rate the following alt text on a scale from 1 to 5 based on its quality in describing the image, assuming the model does not have access to the image: 1 (Very Low Quality): Alt text is vague, irrelevant, misleading, or uses placeholders (e.g., file names). 2 (Low Quality): Alt text is overly simplistic, generic, or provides minimal useful information. 3 (Moderate Quality): Alt text is somewhat descriptive but lacks detail or relevance, with possible redundancy or ambiguity. 4 (High Quality): Alt text is descriptive, clear, concise, and provides sufficient information to understand the image’s content. 5 (Very High Quality): Alt text is highly specific, detailed, and relevant, with a clear description that conveys all key aspects of the image. 2. Subject Classification: Assign a subject/category to the alt text based on its content. Choose from the following categories: Vehicles and Transportation Cooking and Food People and Everyday Life Sports and Recreation Plants and Animals Objects, Materials, and Clothing Brands and Products Geography, Buildings, and Landmarks Tradition, Art, and History Public Figure and Pop-Culture Others 3. Country/Region Classification: Decide if the alt text is closely related to a specific country’s culture. For example, if the alt text says, ”Tokyo Skytree Photo in March with beautiful cherry blossoms”, it’s strongly related to Japan. If the alt text is not specifically about a certain culture or country, you can say ”No specific country”. Even if the alt text is written in their official language, it doesn’t mean the caption is specifically about the country (e.g., a product page caption is often unlikely to be country-specific). Output: Provide the final result in the following structured format: 1. Text Quality Score (1-5): 2. Subject Category: 3. Country/Region: Only generate the final result without any additional descriptions or explanations.
Image Recaption Prompts We randomly select one recaption prompt from the following: PROMPT 1:
Please describe the image in detail in {language}. The image might be related to the country: ”{country}”. The topic might be related to: ”{category}”. The previous short caption of the image is {text}.
PROMPT 2:
Analyze this image and provide a comprehensive description in ”{language}”. Consider that it may be associated with ”{country}” and the theme could be related to ”{category}”. If there is cultural significance, please include it. A brief previous description was: {text}.
PROMPT 3:
In ”{language}”, give a detailed description of what you see in this image. Keep in mind it might be connected to ”{country}” and the subject could be about ”{category}”. If there are culturally relevant details, please include them. An earlier short description stated: {text}.
PROMPT 4:
Examine this image closely and describe its contents in ”{language}” in a more structured way. The image might have a connection to ”{country}” and could be about ”{category}”. A previous concise caption mentioned: {text}.
PROMPT 5:
Using ”{language}”, provide an in-depth and structured description of this image. It may be related to ”{country}” and the topic could be associated with ”{category}”. A prior brief description was given as: {text}.
Instruction Generation Prompt Task: Generate two instruction-response pair based on the visual content of an image. Choose two task types from the list below to guide the rewriting process: Coding & Debugging Information Seeking Creative Writing Critical Reasoning Planning & Strategy Mathematical Thinking Text Revision & Editing Data Analysis Role Playing & Scenarios Brainstorming & Ideation Advice Seeking & Problem-Solving Learning & Understanding Cultural Interpretation Guidelines: Instruction: Select two different task types from the list above. Make sure the instruction prompts an interpretation or analysis directly tied to what can be visually observed in the image, not just general reasoning. The instruction should require a response that uses details from the image. Avoid generic instructions that can be answered without visual information. Response: Provide a very detailed and structured response that reflects a clear understanding of the implied visual information. Offer multiple perspectives, deep analysis, or step-by-step explanations where applicable. Avoid general responses that could be inferred without observing the image. Responses must rely on interpreting the visual content. Content: Instructions should be varied, challenging, and explore different advanced aspects of the visual scene. Responses must showcase a deep understanding of the image’s visual context, using thoughtful insights where applicable. Output: Provide the output in JSON format with three keys: “task_type”, “instruction” and “response”. Ensure the instruction and response do not mention “based on caption” but instead, refer to the image or simply avoid reference to any external description. Do not include additional text or explanations beyond what is required. Provide both the “instruction” and “response” in {language} but “task_type” in English. Caption: {caption}
Appendix C Recaptioning Example from LAION-Cultural
Figure 8: An example from LAION-Cultural illustrating why the filtered informative alt text helps generate a more informative caption. With the high-quality alt text, the model incorporates important details like “President and CEO of The Walt Disney Company standing in front of a model of Shanghai Disneyland” into the generated caption.
Appendix D Datasets used in PangeaBench
To comprehensively assess the capabilities of Pangea across diverse languages, cultures, and task types, we developed PangeaBench. We list the details of each dataset included in the PangeaBench.
D.1 Multimodal Datasets
  • xGQA (Pfeiffer et al., 2022a): A cross-lingual visual question-answering dataset featuring 9,666 questions in eight languages covering five scripts. The dataset includes 300 unique images from Visual Genome (Krishna et al., 2017). xGQA tests the model’s ability to understand and reason about visual content across multiple languages.
  • MaXM (Changpinyo et al., 2022): A VQA dataset in seven languages and five scripts, with questions and answers in the same language. Images are culturally matched to the target language regions. MaXM specifically addresses the challenge of cultural diversity in multimodal understanding.
  • MaRVL (Liu et al., 2021): A Multicultural Reasoning over Vision and Language dataset in five languages and three scripts, featuring 4,914 culturally diverse images matched to respective languages. MaRVL focuses on evaluating models’ ability to reason about culturally diverse visual concepts.
  • XM100 (Thapliyal et al., 2022): We create a subset of 3600 instances (100 instances per language) from the original XM100 dataset, a large multilingual image captioning dataset comprising 36 languages, with 261,375 captions for 100 unique images per language, culturally matched to each language. XM100 evaluates a model’s ability to generate culturally appropriate captions across a wide range of languages. For sampling, we select 100 instances per language, ensuring that all languages share the same set of images for their respective 100 instances. To ensure diversity within our sample, we use Sentence-BERT (Reimers & Gurevych, 2019) to cluster the 3600 English instances from the original dataset into 100 groups, and then select one instance from each group. This method ensures that the sampled instances are as diverse as possible. We evaluate models on this new sample of 3600 instances, which allows for a more time-efficient evaluation while still accurately reflecting the multilingual capabilities of models in diverse contexts.
  • M3Exam (Zhang et al., 2023): A novel benchmark sourced from real and official human exam questions, featuring 12,317 questions in 9 languages across three educational levels. Approximately 23% of the questions require image processing. M3Exam tests the model’s ability to handle complex, multi-step reasoning tasks in an educational context.
  • xMMMU: MMMU contains multimodal questions from college-level materials across six disciplines and 30 subjects. The dataset features 183 subfields and 30 diverse image types, including charts, diagrams, and chemical structures. We sample 300 questions from the original MMMU validation set and translate them using GPT-4o into xx languages. To ensure the quality, we translated each sampled question multiple times and then back-translated it to English. We select the translation with the highest BLEU score. xMMMU evaluates the model’s capacity to understand and reason about specialized academic content across languages and modalities.
D.2 Text-Only Multilingual Datasets
  • TyDiQA (Clark et al., 2020): A question answering dataset covering 11 typologically diverse languages with 204K question-answer pairs. Questions are written by native speakers without seeing the answers, ensuring a realistic information-seeking task. TyDiQA is designed to test linguistic diversity and avoid translation artifacts.
  • FLORES (NLLB Team, 2024): A machine translation benchmark for 200 languages, including many low-resource languages. It consists of 3,001 sentences from 842 web articles, divided into dev, devtest, and test splits. FLORES-200 includes translations from multiple pivot languages and provides script alternatives for some languages, making it a comprehensive test of translation capabilities.
  • MMMLU (OpenAI, 2024a): A human-translated version of MMLU (Hendrycks et al., 2021), covering 57 subjects across STEM, humanities, social sciences, and more. It ranges in difficulty from elementary to advanced professional levels, testing both world knowledge and problem-solving ability in a zero-shot and few-shot setting across multiple languages.
  • MGSM (Shi et al., 2022): Multilingual Grade School Math Benchmark, featuring 250 grade-school math problems translated into 10 languages. Based on GSM8K, it requires multi-step reasoning and tests the model’s ability to solve complex mathematical word problems across languages.
This diverse set of datasets in PangeaBench allows for a comprehensive evaluation of Pangea’s capabilities across various languages, cultures, modalities, and task types, providing a holistic assessment of its performance in multilingual and multimodal contexts.
Appendix E Explanation of xChatBench
Task Category
We first divide into 10 task categories, namely art_explanation, bar_chart_interpretation, defeasible_reasoning, figurative_speech_explanation, iq_test, ocr, graph_interpretation, image_humor_understanding, science_figure_explanation, unusual_images. The task categories are inspired by existing papers that do not use a free-form generation format (Lu et al., 2024; Yue et al., 2024a; Han et al., 2023; Hessel et al., 2023; Kim et al., 2022).
Construction Procedure
To annotate the instances, we mainly follow the procedure of Kim et al. (2024a). Two human annotators first hand-crafted the instances by searching through appropriate images for the task and then hand-crafting each component of the instance. As our motivation for fine-grained evaluation, each instance consists of not only an instruction, reference answer, but also a unique evaluation criteria tailored to each instance (e.g., Does the response effectively explain the humor in the image based on the juxtaposition of a character’s portrayal in different scenarios?) and a description for each score between 1 and 5 (e.g., score4_description: The response understands the juxtaposition and relates it to the humor involving machine learning models, but may miss some nuances or the related aspect of the humor). During the annotation process, we asked the annotators to not copy-and-paste results from LLM services like ChatGPT or directly from the web. Then, we hire four additional annotators to assess the quality of the instances. Each participant to asked to grade if each instance (1) fits into the devised task category, (2) if the quality of the reference answer is good enough, and (3) if the score rubric is suitable to assess the response. We iteratively ask the annotators who made the instances to revise them if the instance does not satisfy all three criteria. The resulting dataset consists of 50 instructions, reference answers, and evaluation criteria with a corresponding score rubric.
Translation Procedure
To assess the multilingual generation capabilities of MLLMs, we translate the hand-crafted 50 instances into 6 different languages, namely Chinese, Hindi, Indonesian, Japanese, Korean, and Spanish. We first use GPT-4o-2024-08-06 to translate the instruction and reference answer of each instance with a naive prompt, “Translate the following sentences into {target_language}. Sentences: {sentences}”. Then, the coauthors who are native speakers of each language reviewed the instances and made adjustments if the translated results were unnatural.
Evaluation Pipeline
Similar to prior works employing LLM-as-a-Judge, we use GPT-4o-2024-08-06 as the judge model and prompt it in a direct assessment manner. As input, the judge model is given the instruction, the model’s response, the reference answer, the evaluation criteria, and the descriptions for each score. As output, the judge generates verbal feedback and an integer score between 1 and 5. For this procedure, we use the prometheus-eval library (Kim et al., 2024b) and employ their default hyperparameter setting for evaluation. Lastly, the final score is acquired by averaging the results across the 50 instances for each language. Note that in the main result and breakdown result tables, we normalize the score from 1-5 to 0-100 by $(\text{score}-1)\times 25$. For the multimodal chat scenarios, we found that many English-centric models tend to respond in English regardless of the query language. This behavior is problematic, as it undermines the fundamental capability of a multilingual model, which should ideally respond in the language of the query. To address this, we implemented a strict evaluation criterion where such responses were penalized and assigned a score of 0. We believe this is crucial, as users may not understand English, and failing to respond in the appropriate language can hinder effective communication and user experience. Thus, for postprocessing, we use langdetect  [^2] https://pypi.org/project/langdetect/ to identify whether the response is written in the given language and change the score to 1 when it is written in a different language, a phenomenon called language hallucination (Xue et al., 2021; Pfeiffer et al., 2023).
Appendix F Qualitative Examples from xChatBench
We show some interesting qualitative examples from xChatBench, including good cases in Figure 9, Figure 10, Figure 11, Figure 12, Figure 13, Figure 14, as well as bad cases in Figure 15, Figure 16. The examples included the scoring rubric, query, response from our Pangea, reference answer, and LLM-as-Judge feedback.
Figure 9: An xChat example from Japanese subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Figure 10: An xChat example from Hindi subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Figure 11: An xChat example from the Korean subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Figure 12: An xChat example from the Indonesian subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Figure 13: An xChat example from Spanish subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Figure 14: An xChat example from the Chinese subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Figure 15: An xChat example from the Chinese subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Figure 16: An xChat example from the Korean subset and our model Pangea’s response. The English part of the instruction, response, and reference answer is additionally added only on the figure to help to understand and was not given nor generated by Pangea.
Appendix G Languages in PangeaIns
Table 5 demonstrates the number of instances for each language that we include in PangeaIns.
Languages en multi am ar bg bn cs de el es
Count 2554.6 4389.5 31.7 162.8 52.7 118.4 4.6 132.2 7.3 126.6
Percentage (%) 36.8 63.2 0.5 2.3 0.8 1.7 0.1 1.9 0.1 1.8
Languages fa fr ga hi id ig it iw ja jv
Count 8.4 193.8 34.7 163.5 178.4 26.1 68.8 168.3 403.3 35.2
Percentage (%) 0.1 2.8 0.5 2.4 2.6 0.4 1.0 2.4 5.8 0.5
Languages ko nl mn ms no pl pt ro ru si
Count 182.5 4.5 37.6 39.2 60.7 8.2 176.5 147.2 211.0 0.6
Percentage (%) 2.6 0.1 0.5 0.6 0.9 0.1 2.5 2.1 3.0 0.1
Languages su sw ta te th tr uk ur vi zh
Count 33.0 132.7 37.1 64.6 160.4 140.7 7.4 156.4 315.5 557.4
Percentage (%) 0.5 1.9 0.5 0.9 2.3 2.0 0.1 2.3 4.5 8.0
Table 5: Language distribution of PangeaIns. We demonstrate the number of instances (in thousands) for each language, and the percentage of instances for each language in PangeaIns
Appendix H Training Examples
H.1 Machine Translated Instructions
We include a few machine-translated training examples from PangeaIns in Figure 17, Figure 18, Figure 19, Figure 20, Figure 21, Figure 22.
H.2 Multicultural Understanding Instructions
We include a few multicultural understanding instructions from PangeaIns in Figure 23, Figure 24, Figure 25, Figure 26.
Figure 17: Translated Task: An example from the Cambrian dataset where it discusses the concept of hourly wages based on a given prompt in Russian.
Figure 18: Multimodal Translated Task: An example from the ALLaVa-LAION dataset where the GPT model answers a prompt in Turkish regarding the maximum time displayed on a digital timer.
Figure 19: Multimodal Translated Task: An example from the ShareGPT-4v dataset where the model describes an image of a wine rack in Thai, detailing its structure and the arrangement of wine bottles.
Figure 20: Text-only Translated Task: An example from the OpenHermes2.5 dataset translated into Japanese.
Figure 21: Text-only Translated Task: An example from the NumininaMath dataset translated into Spanish.
Figure 22: Text-only Translated Task: An example from the Code-Feedback dataset translated into German.
Figure 23: Multicultural Understanding: An example from the LAION-Culture dataset where the GPT model describes and analyzes the Petrobras logo in Portuguese.
Figure 24: Multicultural Understanding: Example from the LAION-Culture dataset where the model interprets the meaning of a Japanese character and its representation in calligraphy, described in Bulgarian.
Figure 25: Multicultural Understanding: LAION-Culture sample in Hebrew.
Figure 26: Multicultural Understanding: LAION-Culture sample in Tamil.
Appendix I Breakdown Results of Different Languages on PangeaBench
I.1 xChat
We show the performance of different models on the xChat benchmark in Table 6.
Models English Multi Spanish Hindi Indonesian Japanese Korean Chinese
Gemini-1.5-Pro 71.0 65.6 66.0 62.0 65.5 68.0 66.5 65.5
GPT4o 67.0 65.1 66.0 64.0 65.0 66.5 67.5 61.5
Llava-1.5-7B 22.5 16.7 22.5 3.5 18.0 23.0 12.0 21.0
Llava-Next-7B 40.5 20.4 33.0 1.5 19.0 25.0 15.0 29.0
Phi-3.5-Vision 38.5 21.1 37.0 11.5 10.5 31.0 12.5 24.0
Cambrian-8B 27.5 15.8 22.5 4.0 20.0 20.0 10.5 18.0
Llava-OV-7B 51.0 33.1 45.5 6.5 42.0 36.5 26.0 42.0
Molmo-7B-D 49.5 34.7 45.0 19.5 36.5 36.0 35.0 46.0
Llama3.2-11B 49.0 31.3 42.5 19.5 45.0 26.0 21.0 43.0
PaliGemma-3B 6.0 3.8 4.5 0.5 6.5 6.5 2.0 3.0
PALO-7B 27.0 16.2 23.0 3.0 19.0 20.0 13.5 18.5
mBLIP mT0-XL 2.5 0.5 0.0 0.0 0.5 2.0 0.5 0.0
mBLIP BLOOMZ-7B 4.0 1.7 2.0 2.5 2.5 0.0 0.0 3.0
Pangea-7B (Ours) 46.0 35.8 43.5 23.5 34.5 39.0 33.5 40.5
Table 6: Comparison of models on the xChat dataset across different languages.
I.2 Multilingual LLaVABench
We show the performance of different models on the Multilingual LLaVABench benchmark in Table 7.
Models English Multi Arabic Bengali Chinese French Hindi Japanese Russian Spanish Urdu
Gemini-1.5-Pro 103.4 106.6 112.9 117.1 104.1 115.5 106.2 118.1 95.7 88.2 101.6
GPT4o 104.6 100.4 98.3 111.9 96.5 101.1 99.7 104.0 88.5 100.9 102.5
Llava-1.5-7B 66.1 40.8 26.4 11.9 50.7 63.8 23.2 70.0 46.5 59.2 15.4
Llava-Next-7B 78.9 50.7 24.9 11.2 72.8 91.4 18.0 70.1 71.8 82.9 13.4
Phi-3.5-Vision 70.8 58.0 50.1 35.1 69.2 86.0 35.9 63.0 67.6 75.6 39.3
Cambrian-8B 78.4 61.8 54.1 35.4 80.9 87.3 44.2 64.4 76.4 90.3 23.3
Llava-OV-7B 89.7 55.3 45.5 33.8 90.0 89.4 35.3 70.3 44.7 75.5 13.3
Molmo-7B-D 95.9 13.8 10.1 4.2 0.3 59.6 5.5 6.0 8.7 29.5 0.0
Llama3.2-11B 93.9 58.2 39.4 48.1 47.2 85.6 67.8 53.7 68.5 77.8 35.3
PaliGemma-3B 32.1 31.9 37.3 38.2 29.1 30.0 35.8 33.4 26.1 32.3 25.1
PALO-7B 68.9 71.2 79.1 54.6 71.5 83.9 61.9 66.6 80.9 74.4 68.2
mBLIP mT0-XL 32.7 28.2 33.7 26.2 3.6 39.8 26.9 26.8 34.1 36.9 26.0
mBLIP BLOOMZ-7B 43.5 41.0 48.1 44.1 30.6 53.3 39.1 29.8 38.1 51.5 34.0
Pangea-7B (Ours) 84.2 89.5 91.0 94.9 94.4 93.8 84.9 92.8 91.2 87.4 75.5
Table 7: Comparison of models on the Multilingual LLaVABench benchmark across different languages.
I.3 CVQA
We show the performance of different models on the CVQA benchmark in Table 8 and Table 9.
Models ar-es br-pt bu-bg ch-es ch-zh co-es ec-es eg-ar et-am et-or
Llava-1.5-7B 37.8 51.1 35.6 42.4 44.4 50.6 48.6 31.5 27.8 31.8
Llava-Next-7B 52.5 62.3 41.5 59.0 51.1 54.8 50.8 33.5 29.5 36.9
Phi-3.5-Vision 54.0 57.2 36.9 57.7 51.1 52.3 50.1 38.4 27.8 32.2
Cambrian-8B 59.6 60.6 42.0 64.5 59.5 57.7 56.1 40.9 27.8 25.7
Llava-OV-7B 64.5 69.7 49.6 67.1 69.1 66.8 65.5 47.8 32.5 41.1
Molmo-7B-D 61.1 69.0 54.9 60.7 66.2 58.5 54.9 56.7 58.1 60.7
Llama3.2-11B 69.1 74.6 64.2 70.5 73.6 69.3 66.9 68.5 68.4 63.1
PaliGemma-3B 48.7 53.9 39.1 53.4 53.7 50.6 45.3 40.4 24.8 28.0
PALO-7B 50.9 56.7 36.7 55.1 45.3 48.5 46.4 28.6 19.2 32.7
mBLIP BLOOMZ-7B 45.3 51.4 30.5 45.3 51.1 46.9 44.8 35.9 23.9 25.7
mBLIP mT0-XL 40.8 44.4 38.0 44.9 39.9 41.9 42.5 31.0 35.9 26.6
Pangea-7B (Ours) 68.3 72.9 53.9 70.5 74.0 64.7 63.5 49.3 36.3 35.5
Models fr-br in-bn in-ta in-te ind-id ind-jv ind-mi ind-sv ir-ir ja-jp
Llava-1.5-7B 29.4 31.1 29.8 28.0 41.7 32.0 32.7 33.5 42.6 37.4
Llava-Next-7B 27.4 31.1 28.8 28.0 42.2 38.7 40.2 35.5 42.6 32.5
Phi-3.5-Vision 29.3 39.0 40.0 36.8 45.0 38.2 38.2 30.8 39.6 39.7
Cambrian-8B 31.6 47.2 38.1 44.0 50.2 43.8 39.4 45.5 47.9 40.9
Llava-OV-7B 34.3 56.3 43.9 46.5 58.0 45.8 45.4 40.5 50.6 49.8
Molmo-7B-D 44.2 61.9 61.2 58.5 52.9 53.9 54.6 55.0 64.4 42.9
Llama3.2-11B 49.4 76.9 80.4 80.5 65.8 60.6 68.9 64.0 76.4 54.2
PaliGemma-3B 29.9 46.2 46.0 43.5 45.4 41.4 39.8 33.0 34.4 43.3
PALO-7B 29.1 37.8 31.2 25.0 41.3 32.3 32.3 32.0 42.9 30.5
mBLIP BLOOMZ-7B 26.7 41.9 40.0 42.0 41.9 35.4 35.1 32.0 29.4 31.0
mBLIP mT0-XL 23.5 36.4 44.2 39.0 37.4 37.4 34.7 31.0 35.3 30.0
Pangea-7B (Ours) 34.6 59.1 51.9 54.5 62.1 49.5 47.8 53.0 56.4 48.3
Models ke-sw ma-my me-es mo-mg ni-ig no-ng pk-ur ph-fi ro-ro ru-ru
Llava-1.5-7B 34.4 42.2 42.4 26.9 34.5 47.5 26.4 43.8 47.0 51.0
Llava-Next-7B 46.2 45.7 51.4 33.3 35.0 56.9 36.6 46.8 52.3 53.5
Phi-3.5-Vision 46.0 45.1 46.3 31.9 33.3 50.0 35.2 41.4 47.4 50.5
Cambrian-8B 50.5 52.1 56.7 34.6 36.0 53.5 48.6 47.3 52.0 61.5
Llava-OV-7B 46.5 55.6 59.4 35.9 33.5 62.5 58.3 56.2 60.3 75.5
Molmo-7B-D 73.3 54.6 53.6 51.9 53.0 54.8 67.1 57.6 63.6 61.5
Llama3.2-11B 79.1 72.1 66.6 54.5 61.5 66.9 78.7 70.0 76.8 74.5
PaliGemma-3B 44.0 44.1 47.4 29.2 32.0 52.2 44.9 39.9 50.3 53.5
PALO-7B 35.9 42.5 44.3 28.8 29.5 49.2 44.4 39.4 46.0 47.0
mBLIP BLOOMZ-7B 37.0 42.5 44.8 28.8 33.0 49.2 47.7 31.5 46.0 34.0
mBLIP mT0-XL 45.1 40.6 44.9 29.2 30.5 42.8 40.3 32.0 43.7 42.0
Pangea-7B (Ours) 64.1 59.7 62.2 42.3 46.0 64.5 66.2 58.6 64.6 74.0
Models rw-ki sg-zh sk-ko sp-es sr-si ur-es macro
Llava-1.5-7B 31.1 44.3 44.5 56.9 24.9 37.8 38.7
Llava-Next-7B 34.5 44.8 43.4 63.5 29.8 41.0 42.6
Phi-3.5-Vision 31.1 43.9 55.2 62.4 28.0 43.3 42.4
Cambrian-8B 31.9 54.7 54.5 70.4 36.4 45.7 47.5
Llava-OV-7B 35.3 70.3 65.2 79.9 31.6 47.3 53.8
Molmo-7B-D 57.4 69.3 65.2 70.1 68.0 50.8 59.4
Llama3.2-11B 57.9 80.7 73.8 81.4 72.4 52.4 70.1
PaliGemma-3B 27.2 48.6 61.0 60.1 31.6 39.4 43.0
PALO-7B 28.9 45.8 44.5 64.8 28.0 39.4 39.3
mBLIP BLOOMZ-7B 29.4 47.6 33.1 56.6 28.0 39.4 36.9
mBLIP mT0-XL 33.2 36.8 38.3 53.5 31.1 39.1 37.6
Pangea-7B (Ours) 35.7 65.6 70.7 72.6 39.1 49.8 57.2
Table 8: Comparison of models on CVQA across different country-language pairs (in local languages). Includes Macro-Acc.
Models ar-es br-pt bu-bg ch-es ch-zh co-es ec-es eg-ar et-am et-or
Llava-1.5-7B 56.2 61.6 52.3 60.2 54.0 55.6 55.5 50.2 51.3 53.3
Llava-Next-7B 53.9 61.3 50.9 59.8 58.8 60.2 52.8 54.7 52.9 58.9
Phi-3.5-Vision 59.2 61.9 54.9 64.1 58.2 59.3 57.5 50.7 54.7 58.4
Cambrian-8B 57.7 66.5 56.1 65.4 64.3 59.3 60.2 56.7 60.3 56.5
Llava-OV-7B 63.0 73.9 59.3 65.8 68.8 65.1 63.3 62.1 59.8 59.3
Molmo-7B-D 57.7 65.8 45.6 63.7 68.5 57.3 55.0 43.8 31.6 38.8
Llama3.2-11B 66.8 72.9 54.4 72.6 72.0 66.4 65.2 56.7 41.9 32.2
PaliGemma-3B 51.7 59.5 49.3 51.7 54.9 54.8 47.2 51.2 52.6 51.4
PALO-7B 50.2 57.0 48.8 53.4 52.1 51.9 53.0 48.3 47.0 52.3
mBLIP mT0-XL 38.1 45.4 39.1 42.7 43.7 41.1 40.9 42.9 34.2 42.1
mBLIP BLOOMZ-7B 46.0 51.4 41.5 44.4 48.9 49.0 45.0 45.3 38.9 46.3
Pangea-7B (Ours) 67.2 72.9 60.1 68.8 67.2 64.7 61.6 59.1 60.7 56.0
Models fr-br in-bn in-ta in-te ind-id ind-jv ind-mi ind-sv ir-ir ja-jp
Llava-1.5-7B 37.3 52.1 61.4 63.5 47.8 50.8 49.0 44.0 61.3 41.9
Llava-Next-7B 37.5 60.8 61.4 60.5 48.5 48.1 51.4 49.0 66.6 40.9
Phi-3.5-Vision 41.7 58.7 60.5 60.0 51.7 45.5 51.4 47.5 62.6 41.4
Cambrian-8B 40.7 68.5 65.6 63.0 55.1 50.2 58.2 56.0 66.6 42.4
Llava-OV-7B 44.2 69.6 72.0 70.5 59.0 55.9 59.4 58.5 76.4 47.3
Molmo-7B-D 29.6 47.9 36.4 41.5 50.5 45.1 43.4 39.5 43.6 44.8
Llama3.2-11B 36.3 62.9 66.4 66.5 63.6 48.8 58.2 54.0 57.4 58.1
PaliGemma-3B 37.3 59.1 66.0 62.5 49.3 48.1 43.4 46.0 58.3 44.8
PALO-7B 36.8 52.4 53.5 56.5 45.1 45.8 44.2 42.0 55.6 37.4
mBLIP mT0-XL 30.4 43.0 46.0 41.0 38.1 39.1 38.6 32.5 37.4 34.0
mBLIP BLOOMZ-7B 34.6 43.4 52.6 49.5 41.0 44.8 38.2 30.5 42.3 36.5
Pangea-7B (Ours) 45.2 67.1 71.0 68.0 60.4 57.2 56.9 56.0 72.7 45.8
Models ke-sw ma-my me-es mo-mg ni-ig no-ng pk-ur ph-fi ro-ro ru-ru
Llava-1.5-7B 68.9 52.1 47.9 45.8 51.0 58.5 63.9 52.7 55.6 59.0
Llava-Next-7B 71.1 54.9 51.1 44.2 53.0 57.2 67.1 56.7 62.6 58.5
Phi-3.5-Vision 72.9 57.1 46.3 50.7 53.0 56.2 60.6 57.6 61.9 58.5
Cambrian-8B 74.4 61.9 56.7 48.7 56.5 60.5 73.1 60.1 66.6 61.5
Llava-OV-7B 79.1 65.1 63.2 52.6 57.5 64.2 75.0 64.0 72.5 72.5
Molmo-7B-D 47.6 51.7 55.1 35.9 36.0 49.2 46.8 43.3 52.0 63.5
Llama3.2-11B 61.5 69.2 64.7 41.0 39.5 65.9 65.7 66.0 75.5 74.5
PaliGemma-3B 59.7 54.9 51.7 43.4 46.0 55.2 67.6 48.8 60.9 56.0
PALO-7B 65.9 49.2 53.4 42.9 49.0 54.5 60.6 52.7 55.0 53.5
mBLIP mT0-XL 50.2 41.6 34.7 33.9 39.5 43.1 45.4 36.9 43.7 41.0
mBLIP BLOOMZ-7B 54.6 45.7 39.3 38.1 45.0 47.2 60.6 36.9 50.3 44.0
Pangea-7B (Ours) 77.2 62.5 61.6 52.9 59.5 64.9 72.2 64.0 71.9 68.5
Models rw-ki sg-zh sk-ko sp-es sr-si ur-es macro
Llava-1.5-7B 51.1 60.8 56.9 66.0 58.7 42.5 54.2
Llava-Next-7B 52.8 62.3 60.0 67.6 59.1 38.7 55.7
Phi-3.5-Vision 52.3 59.4 66.5 66.7 61.3 46.3 56.3
Cambrian-8B 56.2 66.0 63.1 71.7 63.1 47.0 59.7
Llava-OV-7B 55.7 73.6 67.9 80.2 72.9 48.9 65.2
Molmo-7B-D 34.9 66.0 56.9 66.7 31.6 44.8 48.3
Llama3.2-11B 40.4 73.6 73.1 83.3 51.1 56.2 61.2
PaliGemma-3B 44.7 59.4 58.3 61.0 62.2 40.6 52.9
PALO-7B 51.9 56.1 55.9 62.9 54.2 42.2 50.9
mBLIP mT0-XL 38.3 43.9 41.4 51.9 48.0 34.9 40.5
mBLIP BLOOMZ-7B 45.1 53.8 46.9 58.5 46.7 34.0 44.9
Pangea-7B (Ours) 56.6 71.7 66.6 75.2 70.6 52.7 64.4
Table 9: Comparison of models on CVQA across different country-language pairs (in English). Includes Macro-Acc.
I.4 MaRVL
We show the performance of different models on the MaRVL benchmark in Table 10.
Models English Multi Indonesian Swahili Tamil Turkish Chinese
GPT4o 81.8 82.3 81.9 80.8 80.2 86.4 82.1
Gemini-1.5-Pro 76.4 72.0 71.2 67.8 70.0 75.4 75.8
Llava-1.5-7B 56.2 53.7 56.1 49.8 49.7 55.4 57.5
Llava-Next-7B 62.8 50.9 52.2 50.6 50.5 50.4 50.6
Phi-3.5-Vision 72.1 56.5 58.6 51.4 52.0 58.6 61.7
Cambrian-8B 75.4 61.8 64.7 53.6 56.7 65.2 68.9
Llava-OV-7B 72.7 57.5 60.9 51.2 51.9 63.5 60.0
Molmo-7B-D 65.3 54.9 61.1 49.6 49.6 52.2 62.2
Llama3.2-11B 64.5 58.1 62.7 52.4 54.0 61.6 59.5
PaliGemma-3b 56.5 52.2 53.4 49.6 50.5 56.3 51.3
PALO-7B 63.3 54.2 58.3 50.6 51.9 54.9 55.3
mBLIP mT0-XL 67.3 66.7 64.9 64.8 69.7 68.1 65.9
mBLIP BLOOMZ-7B 62.3 58.6 59.1 56.2 60.3 57.7 59.7
Pangea-7B 87.0 79.0 81.3 75.1 69.4 84.8 84.3
Table 10: Comparison of models on the MaRVL dataset across different languages.
I.5 XM100
We show the performance of different models on the XM100 benchmark in Table 11.
Models English Multi Arabic Bengali Czech Danish German Greek
Gemini-1.5-Pro 27.6 19.1 1.7 7.5 25.9 32.8 27.6 5.0
GPT4o 27.7 19.1 15.8 13.5 21.1 25.3 19.3 21.1
Llava-1.5-7B 28.6 1.1 0.0 0.0 2.1 1.0 3.1 0.0
Llava-Next-7B 29.3 9.4 5.6 0.1 12.1 15.7 14.4 4.2
Phi-3.5-Vision 30.2 5.2 0.4 2.4 16.6 16.2 0.0 20.7
Cambrian-8B 20.6 9.9 1.4 6.6 7.4 15.1 15.5 4.4
Llava-OV-7B 30.6 7.0 0.2 0.6 5.2 16.8 14.0 0.4
Molmo-7B-D 22.1 9.1 5.4 7.9 5.7 13.8 12.2 4.2
Llama3.2-11B 27.6 4.5 0.0 0.0 1.5 11.8 4.6 1.2
PaliGemma-3B 18.7 0.8 0.0 0.0 1.1 3.1 2.7 0.0
PALO-7B 30.4 0.8 0.0 0.0 2.0 1.0 2.7 0.0
mBLIP mT0-XL 31.9 3.1 3.2 1.6 3.7 2.1 2.9 3.1
mBLIP BLOOMZ 22.5 10.3 9.5 6.4 11.5 15.9 14.5 10.9
Pangea-7B (Ours) 30.4 14.2 18.1 16.4 16.2 20.7 20.6 11.2
Models Spanish Persian Finnish Filipino French Hebrew Hindi Croatian
Gemini-1.5-Pro 39.5 4.2 29.0 28.7 42.4 4.3 2.2 33.8
GPT4o 28.3 26.6 13.1 26.4 23.1 20.4 17.0 19.4
Llava-1.5-7B 3.7 0.0 0.4 1.1 2.0 0.1 0.0 0.3
Llava-Next-7B 23.6 9.4 5.5 9.3 23.0 2.7 10.2 7.5
Phi-3.5-Vision 20.7 0.0 1.0 1.7 21.2 0.3 0.0 0.5
Cambrian-8B 18.6 9.6 5.1 19.6 18.3 5.8 6.8 7.2
Llava-OV-7B 24.9 3.8 1.5 4.2 22.0 0.0 4.4 7.2
Molmo-7B-D 19.8 11.3 3.1 13.0 19.8 8.3 9.4 6.9
Llama3.2-11B 10.2 0.0 2.4 8.4 12.0 0.0 0.2 0.7
PaliGemma-3B 0.7 0.0 0.1 0.1 0.6 0.0 0.0 1.3
PALO-7B 1.5 0.0 0.4 0.9 2.1 0.0 0.0 0.2
mBLIP mT0-XL 8.3 5.5 1.7 2.8 6.4 4.0 1.8 0.9
mBLIP BLOOMZ 18.9 13.8 4.8 7.7 19.1 7.5 10.1 3.2
Pangea-7B (Ours) 26.2 19.3 3.8 18.9 26.7 18.2 17.4 10.8
Models Hungarian Indonesian Italian Japanese Korean Maori Dutch Norwegian
Gemini-1.5-Pro 37.2 55.4 27.6 1.2 8.2 3.8 27.7 36.7
GPT4o 21.8 28.4 21.0 0.0 11.1 26.8 26.4 24.7
Llava-1.5-7B 3.3 0.9 4.3 0.0 0.0 0.2 2.9 3.7
Llava-Next-7B 9.3 14.7 17.6 4.2 5.2 9.2 23.8 16.3
Phi-3.5-Vision 3.4 3.2 17.5 1.6 0.3 0.2 17.2 14.1
Cambrian-8B 6.6 15.7 15.5 7.2 2.0 3.2 20.3 16.0
Llava-OV-7B 3.6 16.4 12.8 0.6 0.0 1.7 24.7 13.9
Molmo-7B-D 3.5 17.2 17.8 5.2 2.4 7.5 15.7 13.8
Llama3.2-11B 12.7 1.2 16.0 0.0 0.0 9.3 22.0 1.1
PaliGemma-3B 2.0 0.2 1.8 0.0 0.0 4.0 2.6 2.3
PALO-7B 3.4 1.1 3.2 0.0 0.0 0.1 3.5 0.7
mBLIP mT0-XL 2.8 6.0 2.8 0.3 2.1 1.5 3.4 3.1
mBLIP BLOOMZ 11.8 16.0 16.5 0.0 4.5 0.1 18.2 14.5
Pangea-7B (Ours) 7.7 27.9 22.9 2.1 8.1 0.7 26.6 24.9
Models Polish Portuguese Quechua Romanian Russian Swedish Swahili Telugu
Gemini-1.5-Pro 35.5 35.7 0.7 31.2 32.4 37.8 10.7 0.0
GPT4o 22.2 28.0 4.4 19.1 20.7 26.0 20.0 12.5
Llava-1.5-7B 0.8 2.5 0.0 1.6 0.5 2.0 0.1 0.0
Llava-Next-7B 13.5 21.3 0.0 11.5 13.5 16.0 3.2 0.0
Phi-3.5-Vision 1.0 21.0 0.4 3.2 0.7 12.5 0.4 0.0
Cambrian-8B 9.3 17.5 0.0 13.4 11.3 17.9 3.7 2.3
Llava-OV-7B 7.4 24.6 0.0 6.8 5.5 15.0 2.0 0.0
Molmo-7B-D 8.2 16.2 0.6 11.6 12.3 14.1 3.8 0.4
Llama3.2-11B 1.0 18.6 0.0 10.1 0.6 7.4 5.8 0.0
PaliGemma-3B 0.9 1.3 0.1 0.8 0.0 2.0 0.0 0.0
PALO-7B 0.8 1.7 0.0 1.1 0.5 0.9 0.2 0.0
mBLIP mT0-XL 3.5 5.8 0.2 2.3 3.1 3.7 3.8 2.7
mBLIP BLOOMZ 11.8 16.5 0.1 13.7 14.5 14.5 8.4 3.0
Pangea-7B (Ours) 16.2 28.1 0.0 21.4 20.9 19.4 18.7 0.1
Models Thai Turkish Ukrainian Vietnamese Chinese
Gemini-1.5-Pro 0.0 0.9 0.0 0.0 0.9
GPT4o 0.0 17.6 16.9 30.9 0.4
Llava-1.5-7B 0.0 0.0 0.0 0.0 0.0
Llava-Next-7B 0.0 0.0 0.3 0.0 6.3
Phi-3.5-Vision 0.5 1.9 0.0 2.2 0.0
Cambrian-8B 0.4 9.3 5.9 17.8 11.3
Llava-OV-7B 0.0 0.0 0.0 0.0 2.9
Molmo-7B-D 0.0 0.0 0.0 0.0 0.0
Llama3.2-11B 0.0 0.0 0.0 0.0 2.9
PaliGemma-3B 0.5 0.0 0.0 0.2 0.0
PALO-7B 0.2 0.0 0.0 0.1 0.0
mBLIP mT0-XL 0.0 3.9 2.0 7.1 0.0
mBLIP BLOOMZ 0.5 1.9 0.0 2.2 0.0
Pangea-7B (Ours) 0.0 0.0 0.3 0.0 4.9
Table 11: Comparison of models on the XM100 dataset across different languages.
I.6 xGQA
We show the performance of different models on the xGQA benchmark in Table 12.
Models English Multi Bengali German Indonesian Korean Portuguese Russian Chinese
Gemini-1.5-Pro 54.2 48.7 49.4 50.2 48.6 46.4 51.2 44.8 50.2
GPT4o 55.8 51.0 49.4 52.6 50.4 51.0 52.2 50.0 51.4
Llava-1.5-7B 62.0 30.7 15.6 28.4 33.4 38.2 27.5 33.1 38.4
Llava-Next-7B 64.8 37.8 11.5 41.5 37.3 42.5 39.8 43.5 48.2
Phi-3.5-Vision 64.7 38.4 7.7 51.4 36.0 36.3 49.6 46.2 41.4
Cambrian-8B 64.6 39.8 32.3 44.6 36.0 43.6 41.6 44.2 36.2
Llava-OV-7B 64.4 48.2 41.8 49.2 48.8 45.3 52.4 54.0 45.9
Molmo-7B-D 51.5 43.0 25.6 45.9 44.9 44.2 46.5 45.6 48.1
Llama3.2-11B 55.6 45.4 42.9 46.7 46.2 44.5 46.5 44.7 46.1
PaliGemma-3B 59.7 30.5 13.3 44.5 21.3 22.8 34.7 35.8 41.2
PALO-7B 60.5 37.8 42.2 39.1 36.8 41.7 31.7 27.0 46.5
mBLIP mT0-XL 44.2 39.9 39.1 41.1 39.1 39.7 40.7 40.2 39.4
mBLIP BLOOMZ-7B 43.3 36.9 37.7 36.3 39.3 28.5 40.7 36.6 39.1
Pangea-7B (Ours) 64.7 60.2 58.9 61.6 60.1 58.9 61.8 60.4 59.6
Table 12: Comparison of models on the xGQA dataset across different languages
I.7 MAXM
We show the performance of different models on the MAXM benchmark in Table 13.
Models English Multi French Hindi Hebrew Romanian Thai Chinese
Gemini-1.5-Pro 56.4 63.5 60.2 66.5 65.7 57.4 73.9 57.4
GPT4o 60.7 65.4 59.8 68.8 70.0 61.3 76.5 56.3
Llava-1.5-7B 49.8 20.4 32.2 17.3 12.9 15.1 17.2 27.8
Llava-Next-7B 54.9 21.4 33.7 16.2 10.7 15.5 18.3 33.9
Phi-3.5-Vision 55.3 25.0 38.3 31.9 17.5 10.9 24.3 27.4
Cambrian-8B 55.3 28.7 41.7 23.8 17.1 32.0 25.7 31.8
Llava-OV-7B 54.9 34.8 37.9 31.9 17.8 30.2 53.0 37.9
Molmo-7B-D 52.9 37.5 45.5 33.5 30.7 28.9 46.3 40.4
Llama3.2-11B 55.3 43.9 48.1 50.4 41.8 36.6 56.7 30.0
PaliGemma-3B 47.9 19.9 8.0 36.5 19.3 13.4 31.3 10.8
PALO-7B 51.4 16.3 33.7 15.8 12.1 11.3 14.6 10.5
mBLIP mT0-XL 44.7 36.8 36.0 42.7 28.9 30.3 56.3 26.4
mBLIP BLOOMZ-7B 44.7 24.8 33.0 47.3 8.9 16.9 9.7 33.2
Pangea-7B (Ours) 55.3 53.3 43.6 53.5 59.3 45.8 67.2 50.2
Table 13: Comparison of models on the MAXM dataset across different languages.
I.8 xMMMU
We show the performance of different models on the xMMMU benchmark in Table 14.
Models English Multi Arabic French Hindi Indonesian Japanese Portuguese
Gemini-1.5-Pro (0801) 65.8 57.7 57.7 58.1 55.5 60.2 55.0 59.6
GPT4o (0513) 69.1 58.3 56.7 58.1 58.1 59.9 58.0 58.9
Llava-1.5-7B 36.2 31.5 29.5 34.9 27.5 31.6 32.0 33.7
Llava-Next-7B 36.7 34.3 30.5 35.6 30.9 37.0 34.9 37.0
Phi-3.5-Vision 42.6 38.8 35.6 44.0 30.9 36.7 37.9 47.8
Cambrian-8B 41.8 33.2 32.6 34.6 30.9 31.3 33.5 36.0
Llava-OV-7B 46.3 41.0 41.6 43.0 34.7 43.4 40.1 43.4
Molmo-7B-D 42.9 40.4 40.6 42.6 32.6 40.7 43.9 42.1
Llama3.2-11B 39.2 34.0 33.6 39.6 32.3 36.7 29.0 33.0
PaliGemma-3B 26.3 25.2 29.2 23.8 21.6 24.2 24.5 27.6
PALO-7B 33.1 30.5 30.5 33.2 28.9 34.0 27.1 33.3
mBLIP mT0-XL 29.3 30.4 30.2 33.2 28.2 26.9 31.6 32.3
mBLIP BLOOMZ-7B 29.2 30.8 28.5 33.9 27.8 33.3 31.6 29.6
Pangea-7B (Ours) 45.7 43.7 42.3 45.3 41.6 46.5 40.5 46.1
Table 14: Comparison of models on the xMMMU dataset across different languages.
I.9 M3Exam
We show the performance of different models on the M3Exam benchmark in Table 15.
Models English Multi Afrikaans Chinese Italian Portuguese Thai Vietnamese
Gemini-1.5-Pro 77.4 64.7 80.4 74.1 76.3 61.8 49.9 46.0
GPT4o 68.0 61.0 73.0 68.0 67.0 58.0 52.0 48.3
Llava-1.5-7B 32.3 29.0 28.2 24.3 40.1 28.2 23.7 29.3
Llava-Next-7B 36.5 28.4 28.2 25.4 37.8 27.0 23.7 28.4
Phi-3.5-Vision 55.8 37.2 44.2 40.8 51.4 40.3 25.2 21.6
Cambrian-8B 34.7 33.4 36.8 34.2 45.2 30.3 28.9 25.0
Llava-OV-7B 60.4 45.8 50.3 58.0 57.2 43.8 30.9 34.5
Molmo-7B-D 57.1 39.1 35.6 56.4 49.4 40.2 27.4 25.9
Llama3.2-11B 51.8 36.6 42.3 46.4 45.8 28.4 26.4 30.2
PaliGemma-3B 36.0 25.6 26.4 24.7 32.2 24.3 27.2 19.0
PALO-7B 30.8 27.8 31.9 22.1 36.9 32.3 22.7 20.7
mBLIP mT0-XL 22.8 25.0 16.0 25.6 33.7 21.2 22.4 31.0
mBLIP BLOOMZ-7B 30.3 29.5 28.2 29.8 37.3 28.3 22.9 30.2
Pangea-7B (Ours) 61.4 42.1 52.1 49.2 54.9 43.3 32.9 19.8
Table 15: Comparison of models on the M3Exam dataset across different languages.
I.10 TyDiQA
We show the performance of different models on the TyDiQA benchmark in Table 16.
Models English Multi Arabic Bengali Finnish Indonesian Korean Russian Swahili Telugu
Vicuna-1.5-7B 59.7 52.7 32.3 68.1 63.0 72.6 58.8 57.6 51.3 18.1
Qwen2-7B-Instruct 72.2 71.2 67.6 75.9 67.1 78.0 64.9 67.2 75.3 73.8
Llava-1.5-7B 66.8 52.8 61.8 33.4 60.2 72.8 63.3 55.0 55.0 20.6
Llava-Next-7B 68.3 52.1 64.5 24.9 63.0 74.3 61.9 58.4 53.1 17.0
Phi-3.5-Vision 75.9 51.3 63.1 24.8 57.3 70.6 60.2 57.5 48.7 28.3
PALO-7B 69.4 50.8 60.9 46.0 61.8 70.6 56.8 56.7 42.5 10.8
Pangea-7B (Ours) 73.7 66.0 55.5 65.3 66.3 74.5 69.4 60.1 76.6 60.0
Table 16: Comparison of models on the TyDiQA dataset across different languages.
I.11 XStoryCloze
We show the performance of different models on the XStoryCloze benchmark in Table 17.
Models English Multi Arabic Spanish Basque Hindi Ind. Burmese Russian Swahili Telugu Chinese
Vicuna-1.5-7B 78.1 57.4 52.7 69.4 50.8 54.5 61.0 48.4 66.5 52.1 54.5 63.5
Qwen2-7B-Instruct 80.3 61.9 64.0 71.6 51.6 59.6 68.5 50.7 72.7 53.2 55.3 72.1
Llava-1.5-7B 79.1 57.6 52.7 69.2 50.9 54.9 62.6 49.0 65.9 51.7 55.8 63.9
Llava-Next-7B 79.1 57.1 51.7 68.8 50.3 54.5 62.0 46.7 65.5 52.1 55.2 63.8
Phi-3.5-Vision 77.9 54.8 53.7 67.2 50.4 54.9 51.7 47.8 61.3 49.3 52.5 59.5
PALO-7B 77.4 57.2 56.5 68.4 49.8 58.6 58.5 47.4 65.6 51.2 53.1 62.8
Pangea-7B (Ours) 79.1 61.2 60.5 67.8 50.0 61.8 66.4 48.7 69.4 58.9 60.4 68.2
Table 17: Comparison of models on the XStoryCloze dataset across different languages.
I.12 MGSM
We show the performance of different models on the MGSM benchmark in Table 18.
Models English Multi Bengali German Spanish French Japanese Russian Swahili Telugu Thai Chinese
Vicuna-1.5-7B 17.6 6.4 0.0 14.4 9.6 14.4 2.8 10.8 3.6 0.0 2.0 14.8
Qwen2-7B-Instruct 48.8 40.4 0.0 67.2 67.6 68.8 11.2 71.2 10.8 2.4 45.6 59.2
Llava-1.5-7B 14.8 7.6 0.0 15.2 10.8 18.0 2.8 11.2 0.4 0.0 1.6 15.6
Llava-Next-7B 15.6 7.5 0.0 13.6 13.2 16.0 1.6 12.8 2.0 0.0 1.6 14.0
Phi-3.5-Vision 59.2 33.1 0.0 64.0 59.6 58.0 20.0 54.0 4.0 0.0 18.8 52.4
PALO-7B 13.6 5.8 0.0 11.6 9.6 13.2 1.6 8.8 0.4 0.0 0.0 12.4
Pangea-7B (Ours) 82.0 47.4 0.0 68.4 74.8 63.2 22.0 68.0 54.0 5.6 49.6 68.0
Table 18: Comparison of models on the MGSM dataset across different languages.
I.13 MMMLU
We show the performance of different models on the MMMLU benchmark in Table 19.
Models English Multi Arabic Bengali Portuguese Chinese French German
Vicuna-1.5-7B 49.5 34.7 30.3 28.5 39.6 36.9 40.4 39.8
Qwen2-7B-Instruct 70.1 53.1 51.0 43.4 60.7 63.8 61.5 57.7
Llava-1.5-7B 50.2 34.9 29.7 28.5 40.3 36.8 40.1 39.8
Llava-Next-7B 52.1 35.6 30.0 28.8 40.7 37.3 41.4 41.4
Phi-3.5-Vision 62.0 39.1 34.9 27.9 47.6 41.5 49.2 45.8
PALO-7B 46.7 32.6 30.3 29.5 36.0 34.2 36.9 35.8
Pangea-7B (Ours) 68.4 52.2 49.3 44.4 58.9 60.5 58.9 56.7
Models Hindi Indonesian Italian Japanese Korean Spanish Swahili Yoruba
Vicuna-1.5-7B 29.8 36.5 39.5 35.9 34.1 40.3 27.9 26.8
Qwen2-7B-Instruct 45.7 57.1 60.8 58.0 54.6 61.9 36.0 31.8
Llava-1.5-7B 29.2 37.1 41.0 35.1 34.1 41.6 28.0 27.3
Llava-Next-7B 29.6 37.5 41.2 36.0 34.2 42.7 28.5 28.7
Phi-3.5-Vision 32.9 38.3 47.0 40.0 36.6 49.6 28.9 27.8
PALO-7B 29.6 33.7 36.4 32.7 30.6 37.0 26.4 27.1
Pangea-7B (Ours) 45.7 55.4 58.8 55.3 52.7 59.7 42.8 31.3
Table 19: Comparison of models on the MMMLU dataset across different languages.
Appendix J A Preliminary Exploration of Constructing Multilingual OCR Instructions
Optical Character Recognition (OCR) is a critical capability for multimodal LLMs, enabling them to interpret and process textual information embedded within images. However, most existing OCR training datasets are predominantly English-centric, which limits the models’ performance in non-English contexts. To address this gap, we have curated a comprehensive set of 500K multilingual OCR training samples from web user interfaces, spanning 10 languages, with 50K examples per language, sourced from web user interfaces. Webpages naturally serve as image-rich environments containing abundant text, and by capturing screenshots of websites from various countries in different languages, we were able to gather a substantial number of OCR images.
We utilize URLs from the CC-News-Multilingual [^3] https://huggingface.co/datasets/intfloat/multilingual_cc_newsdataset (Hamborg et al., 2017) to obtain a diverse set of multilingual web pages. Using Playwright [^4] https://github.com/microsoft/playwright, we render each website and automatically capture screenshots under various device settings and resolutions to achieve a wide range of image dimensions and aspect ratios. Each screenshot includes a red bounding box that highlights a specific element targeted for OCR extraction. We focus on ten languages for this dataset: English, Chinese, Japanese, Korean, Indonesian, Hindi, Spanish, French, Portuguese, and Arabic. We totally have 1M samples (50K for each language).
Figure 27: A preliminary exploration of multilingual OCR.
We employed the same model architecture as Pangea but trained it exclusively on these OCR images, reserving a portion of the data as a test set. As shown in Figure 27 , the results indicate that improving multilingual OCR performance is feasible with an increase in training samples. However, the OCR accuracy for non-Latin scripts (e.g., Chinese, Japanese, and Korean) remains lower than for Latin-based languages. Looking ahead, we aim to further expand the multilingual OCR training dataset to include more languages and integrate this data into PangeaIns.